# Model-Based Diffusion Sampling for Predictive Control in Offline Decision Making

### v1

@@ Appendix D Inpainting vs Conditioning
Most prior trajectory diffusion works (e.g., janner2022planning) adopt a U-Net architecture that diffuses the entire sequence $(x_{0:T-1},u_{0:T-1})$ and incorporates the initial state $x_{0}$ via an inpainting scheme. In contrast, we propose to inject $x_{0}$ directly through FiLM layers, (perez2018film), while diffusing $(x_{1:T},u_{0:T-1})$. This design ensures that the observed initial state is encoded consistently across the diffusion process without requiring partial trajectory masking. To evaluate the effect of this change, we conduct an ablation on D4RL Hopper tasks. As shown in Table 8, conditioning through FiLM provides a consistent improvement over inpainting across all datasets, suggesting that our conditioning scheme is an effective way to incorporate initial state into trajectory diffusion models.
@@ Appendix G Computation Budget, Replanning Experiment
We analyze the runtime characteristics of our compositional diffusion procedure in the D4RL hopper-medium-expert-v2 environment. After training both the planner and dynamics diffusion models, we generate trajectories according to Algorithm 1. Naively, each environment step requires running a full reverse diffusion chain, which can be computationally expensive. To accelerate planning, we adopt a warm-start strategy: the generated trajectory from the previous step is partially diffused forward for a fixed number of steps, after which the same number of reverse diffusion steps are applied to obtain a new trajectory as proposed in (janner2022planning). Owing to the improved dynamic feasibility of our generated sequences, the warm-started trajectory remains close to the optimal continuation, since the observed environment state is typically very close to the predicted next state. As shown in Figure 15, the number of denoising steps can be reduced substantially with little loss in performance: using only $10$ steps yields an average return of $94.9$ with $58$ ms per action, while beyond $20$ steps performance is comparable to running the full diffusion chain.
@@ Appendix I Comparison of Alternation and Score Combination Methods
We study the effect of alternating versus combined score updates, motivated by Eq. equation 7. In the combined setting, we directly do a convex combination of the planner and dynamics scores over state dimensions at each diffusion step and perform a single denoising update, rather than alternating between the two models. We evaluate both approaches on the D4RL hopper environments. For combined score we do a grid search for weighting parameter and report the best result.
@@
A natural question is whether the denoiser should mirror the forward-time causality of the underlying dynamics or whether such a restriction limits its modeling capacity. Motivated by this, and following observations in chen2024diffusion, we examine the effect of enforcing temporal causality in our denoising networks. We re-implement both the planner and dynamics models using causal U-Nets in the WaveNet style rethage2018wavenet and evaluate them on D4RL medium-replay tasks.
As reported in Table 14, causal architectures lead to a slight drop in performance. Although system dynamics are inherently causal, the optimal denoiser in a diffusion model need not be: score estimation at each timestep is a smoothing operation that benefits from future context wiener1964extrapolation, and similar observations have been made in diffusion models for audio and speech kong2020diffwave. Our results align with this: view—strict causality restricts receptive fields and degrades the quality of the learned score, whereas acausal models exploit full-context information during denoising.
@@
We evaluate the sensitivity of MPDiffuser to two key sampling hyperparameters on FetchPickAndPlace: the classifier-free guidance scale $w$, which controls the strength of task conditioning during denoising, and the number of sampled trajectories used by the ranker. As shown in Tables 16 and 17, performance remains stable over a broad range of guidance strengths, with a mild peak around $w\in[1.5,2.5]$. Increasing the number of samples for ranking yields improvements up to roughly 8 samples, after which returns saturate.
@@
Let $p_{\mathrm{pl}}(\tau\mid x_{0})$ denote the induced trajectory distribution obtained by running a fixed (e.g., DDIM) sampling procedure from the learned score/denoiser, conditioned on the initial state $x_{0}$. Intuitively, $p_{\mathrm{pl}}$ concentrates on trajectories that resemble the dataset and thus capture task structure, multimodality, and preferences present in demonstrations.
@@
The constraint preserves fidelity to the planner—retaining its task‑relevant structure and sample quality—while the objective steers probability mass toward trajectories that are more probable under the dynamics (i.e., higher $p_{\mathrm{dyn}}$). In this sense, equation 14 is a projection of $p_{\mathrm{pl}}$ onto the set of dynamics‑consistent distributions within a KL ball, yielding a principled balance between planner fidelity and dynamics feasibility.
The constrained projection equation 14 can be handled via a Lagrangian relaxation, leading to the unconstrained objective
@@
By Gibbs’ variational principle cover1999elements, the unique minimizer of equation 15 is given by an exponential tilting of the planner distribution:
@@
Directly characterizing $q^{\ast}$ is difficult in practice: we do not have an explicit form for the planner distribution $p_{\mathrm{pl}}$ nor for the dynamics distribution $p_{\mathrm{dyn}}$, and thus cannot evaluate or draw samples from their product-of-experts combination. An alternative is to appeal to the diffusion framework, where one can sample from a target distribution by following a discrete approximation of its probability–flow dynamics. At diffusion step $k$, DDIM update song2020denoising takes the form:
@@
This alternating procedure can be interpreted through the lens of fractional-step or operator-splitting methods (hairer2006geometric; trotter1959product; strang1968construction). When a system evolves under two interacting vector fields—here represented by the planner and dynamics scores—alternating short integration steps under each component provides a first-order Lie–Trotter approximation to the joint flow. For sufficiently small step size, the global discretization error of the Lie–Trotter splitting decays linearly with step size. Hence, as the step size decreases, the alternating process converges to the true combined flow $s_{q^{\ast}}$.
@@
For our diffusion models, we use a planning horizon of $H=16$ with $K=10$ denoising steps. These optimizations together enable real-time operation on the Go2, yielding dynamically feasible trajectories at control rates sufficient for agile locomotion.
Appendix P Use of Large Language Models
We employ LLMs for three main purposes: (1) refining text by condensing content and correcting grammar, (2) suggesting related work that might otherwise be overlooked, and (3) assisting with figure and table formatting. In addition, we use LLM-enabled IDEs to support programming tasks.

### v2

@@ Appendix D Inpainting vs Conditioning
Most prior trajectory diffusion works (e.g., Janner et al. (2022)) adopt a U-Net architecture that diffuses the entire sequence $(x_{0:T-1},u_{0:T-1})$ and incorporates the initial state $x_{0}$ via an inpainting scheme. In contrast, we propose to inject $x_{0}$ directly through FiLM layers, (Perez et al., 2018), while diffusing $(x_{1:T},u_{0:T-1})$. This design ensures that the observed initial state is encoded consistently across the diffusion process without requiring partial trajectory masking. To evaluate the effect of this change, we conduct an ablation on D4RL Hopper tasks. As shown in Table 9, conditioning through FiLM provides a consistent improvement over inpainting across all datasets, suggesting that our conditioning scheme is an effective way to incorporate initial state into trajectory diffusion models.
@@ Appendix G Computation Budget, Replanning Experiment
We analyze the runtime characteristics of our compositional diffusion procedure in the D4RL hopper-medium-expert-v2 environment. After training both the planner and dynamics diffusion models, we generate trajectories according to Algorithm 1. Naively, each environment step requires running a full reverse diffusion chain, which can be computationally expensive. To accelerate planning, we adopt a warm-start strategy: the generated trajectory from the previous step is partially diffused forward for a fixed number of steps, after which the same number of reverse diffusion steps are applied to obtain a new trajectory as proposed in (Janner et al., 2022). Owing to the improved dynamic feasibility of our generated sequences, the warm-started trajectory remains close to the optimal continuation, since the observed environment state is typically very close to the predicted next state. As shown in Figure 13, the number of denoising steps can be reduced substantially with little loss in performance: using only $10$ steps yields an average return of $94.9$ with $58$ ms per action, while beyond $20$ steps performance is comparable to running the full diffusion chain.
@@ Appendix I Comparison of Alternation and Score Combination Methods
We study the effect of alternating versus combined score updates, motivated by Eq. (2). In the combined setting, we directly do a convex combination of the planner and dynamics scores over state dimensions at each diffusion step and perform a single denoising update, rather than alternating between the two models. We evaluate both approaches on the D4RL hopper environments. For combined score we do a grid search for weighting parameter and report the best result.
@@
A natural question is whether the denoiser should mirror the forward-time causality of the underlying dynamics or whether such a restriction limits its modeling capacity. Motivated by this, and following observations in Chen et al. (2024), we examine the effect of enforcing temporal causality in our denoising networks. We re-implement both the planner and dynamics models using causal U-Nets in the WaveNet style (Rethage et al., 2018) and evaluate them on D4RL medium-replay tasks.
As reported in Table 15, causal architectures lead to a slight drop in performance. Although system dynamics are inherently causal, the optimal denoiser in a diffusion model need not be: score estimation at each timestep is a smoothing operation that benefits from future context (Wiener, 1964), and similar observations have been made in diffusion models for audio and speech (Kong et al., 2020). Our results align with this: view—strict causality restricts receptive fields and degrades the quality of the learned score, whereas acausal models exploit full-context information during denoising.
@@
We evaluate the sensitivity of MPDiffuser to two key sampling hyperparameters on FetchPickAndPlace: the classifier-free guidance scale $w$, which controls the strength of task conditioning during denoising, and the number of sampled trajectories used by the ranker. As shown in Tables 17 and 18, performance remains stable over a broad range of guidance strengths, with a mild peak around $w\in[1.5,2.5]$. Increasing the number of samples for ranking yields improvements up to roughly 8 samples, after which returns saturate.
@@
Let $p_{\mathrm{pl}}(\tau\mid x_{0})$ denote the induced trajectory distribution obtained by running a fixed (e.g., DDIM) sampling procedure from the learned score/denoiser, conditioned on the initial state $x_{0}$. Intuitively, $p_{\mathrm{pl}}$ concentrates on trajectories that resemble the dataset and thus capture task structure, and preferences present in demonstrations.
@@
The constraint preserves fidelity to the planner—retaining its task‑relevant structure and sample quality—while the objective steers probability mass toward trajectories that are more probable under the dynamics (i.e., higher $p_{\mathrm{dyn}}$). In this sense, (9) is a projection of $p_{\mathrm{pl}}$ onto the set of dynamics‑consistent distributions within a KL ball, yielding a principled balance between planner fidelity and dynamics feasibility.
The constrained projection (9) can be handled via a Lagrangian relaxation, leading to the unconstrained objective
@@
By Gibbs’ variational principle Cover (1999), the unique minimizer of (10) is given by an exponential tilting of the planner distribution:
@@
Directly characterizing $q^{\ast}$ is difficult in practice: we do not have an explicit form for the planner distribution $p_{\mathrm{pl}}$ nor for the dynamics distribution $p_{\mathrm{dyn}}$, and thus cannot evaluate or draw samples from their product-of-experts combination. An alternative is to appeal to the diffusion framework, where one can sample from a target distribution by following a discrete approximation of its probability–flow dynamics. At diffusion step $k$, DDIM update Song et al. (2021) takes the form:
@@
This alternating procedure can be interpreted through the lens of fractional-step or operator-splitting (Hairer et al., 2006; Trotter, 1959; Strang, 1968). When a system evolves under two interacting vector fields—here represented by the planner and dynamics scores—alternating short integration steps under each component provides a first-order Lie–Trotter approximation to the joint flow. For sufficiently small step size, the global discretization error of the Lie–Trotter splitting decays linearly with step size. Hence, as the step size decreases, the alternating process converges to the true combined flow $s_{q^{\ast}}$.
@@
For our diffusion models, we use a planning horizon of $H=16$ with $K=10$ denoising steps. These optimizations together enable real-time operation on the Go2, allowing control rates sufficient for agile locomotion.